OpenStudy (lgbasallote):

Perform the multiplication:

OpenStudy (lgbasallote):

(xD - 1)D

OpenStudy (lgbasallote):

Note: this is differential equations NOT algebra

OpenStudy (anonymous):

0

OpenStudy (lgbasallote):

lol

OpenStudy (amistre64):

is it rdrr ?

OpenStudy (lgbasallote):

i dont know. how did you do it?

OpenStudy (amistre64):

rdrr is a mnuemonic for har dee har har

OpenStudy (lgbasallote):

lol. seriously do you know how to do this?

OpenStudy (amistre64):

im not sure what you mean by multiply differential

OpenStudy (lgbasallote):

oh :( it's operations on higher order linear DE thingy

OpenStudy (amistre64):

\[\frac{dy}{dx}\frac{dy}{dx}=(\frac{dy}{dx})^2\]

OpenStudy (amistre64):

seems beyond my current knowledge :/

OpenStudy (lgbasallote):

aww :( all i know is it's not equal to D(xD - 1)..other than that im stuck

OpenStudy (apoorvk):

This is a TROLL-QUESTION!!!!!! did you notice that ' xD ' there?? :O

OpenStudy (lgbasallote):

loll

OpenStudy (lgbasallote):

D means differential operator :p x is a variable

mathslover (mathslover):

(xD-1)(D) =\(xD^2-D\) .. good question :D

OpenStudy (lgbasallote):

lol it's not that simple :p

OpenStudy (anonymous):

D is a differential operator that is : \(D = \frac{d}{dx}\).. xD can be written as : \[x \frac{d}{dx} \implies = \frac{dx}{dx} = 1\] Taken the x inside.. So, (xD - 1)D (1 - 1)D = 0 I think so..

OpenStudy (lgbasallote):

so what would be D(xD - 1)?

OpenStudy (anonymous):

Solve inner one first.. xD you can write it as Dx.. Dx = 1 So, \[\frac{d}{dx}(1 - 1) = 0\]

OpenStudy (lgbasallote):

D(xD - 1) shouldnt be equal to (xD - 1)D

OpenStudy (lgbasallote):

so which one is wrong in your solution?

OpenStudy (anonymous):

See, I think (xD - 1)D is wrong actually. This should be D(xD - 1) See the difference: \[(x \frac{d}{dx} - 1) \frac{d}{dx}\] Does it make sense?? Or, \[\frac{d}{dx}(x \frac{d}{dx} - 1)\] This also does not make sense but as variable x can be taken inside so, \[\frac{d}{dx}( \frac{dx}{dx} - 1)\] This makes sense.

OpenStudy (lgbasallote):

uhmm what's your point?

OpenStudy (lgbasallote):

i really dont see what you mean

OpenStudy (anonymous):

D(xD - 1) is right in my point..

OpenStudy (lgbasallote):

so (xD - 1)D is?

OpenStudy (anonymous):

See, you can take (xD - 1) inside; So it will become: D(xD - 1)

OpenStudy (lgbasallote):

\[D(xD - 1) \ne (xD - 1)D\] according to that rule thingy... i dont get what you mean @waterineyes

OpenStudy (anonymous):

Call @UnkleRhaukus , he will give some opinions on this..

OpenStudy (experimentx):

D and xD - 1 do not commute

OpenStudy (experimentx):

D(xD - 1) = D + xD^2 (xD-1)D = xD^2 - D

OpenStudy (lgbasallote):

hmm so it becomes \[x \frac{d^2 y}{dx^2} - \frac{dy}{dx}?\]

OpenStudy (experimentx):

if D is defined as d/dx then ... that is the last one. all right gotta go for a while.

OpenStudy (lgbasallote):

aww..i need to know how you got it :(

OpenStudy (experimentx):

D(xD - 1) = D(xD) - D(1) = D(xD) <-- chain rule =>D(x) D + x DD = D + x D^2 (xD - 1)D = xD(D) - 1(D) = xD^2 - D

OpenStudy (experimentx):

D(1) = 0

OpenStudy (lgbasallote):

OHHHH i see

OpenStudy (unklerhaukus):

\[\qquad(x\text D - 1)\text D=x\text D^2-\text D\] \[\left( x\frac{\text d}{\text dx} -1\right)\frac{\text d}{\text dx}=\left( x\frac{\text d^2}{\text dx^2} -\frac{\text d}{\text dx}\right)\]

OpenStudy (lgbasallote):

how do you make the parentheses big? using array?

OpenStudy (unklerhaukus):

\left( \right)